# A Comprehensive Survey on Deep Multimodal Learning with Missing Modality

### v2

@@ 2.3.4. Multimodal Large Language Model
With the rise of LLMs like ChatGPT (Bubeck et al., 2023), their impressive generalization capabilities across tasks have demonstrated the transformative power. However, human understanding extends beyond language, incorporating modalities like vision and audio. This has led researchers to explore MLLMs, designed to handle diverse user inputs across modalities, including cases with missing modalities, leveraging the flexibility of Transformers. In current MLLM architectures, LLMs act as feature processors, integrating feature tokens from different modality-specific encoders and passing the output to task-/modality-specific decoders. This enables LLM to not only capture rich inter-modal dependencies, but also naturally carry the ability to handle any number of modalities, that is, the ability to solve the missing modality problem. Most MLLMs employ transformer-based modality encoders, such as CLIP, ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023), which encode multimodal inputs into a unified representation space. Examples include BLIP-2 (Li et al., 2023b), which bridges the modality gap using a lightweight Querying Transformer, and LLaVA (Liu et al., 2024a), which enhances visual-language understanding with GPT-4-generated instruction data (Bubeck et al., 2023). These models are optimized for tasks like Visual Question Answering, Dialogue, and Captioning. Recent advancements extend output generation to multiple modalities, such as images. AnyGPT (Zhan et al., 2024) and NExT-GPT (Wu et al., 2023a) unify modalities like text, speech, and images using discrete representations and multimodal adaptors, enabling seamless multimodal interaction. CoDi (Tang et al., 2024) introduces Composable Multimodal Conditioning, allowing arbitrary modality generation through weighted feature summation. Other notable MLLMs include Qwen-VL (Bai et al., 2023), AnyMAL (Moon et al., 2023), and X-LLM (Chen et al., 2023b), showcasing the growing potential of MLLMs in diverse applications. Although MLLMs can flexibly handle any number of modalities, they have disadvantages of inconsistent multi-modal positional encoding, training difficulty, and high GPU resource requirements. Additionally, no specific MLLM benchmarks on missing modality problems have been found.
2.4. Model Selection
@@
Model selection methods aim to use one or more selected models for downstream tasks, while also enhancing robustness and performance. These methods can be categorized into ensemble, dedicated, and discrete scheduler methods. Ensemble methods combine predictions from multiple selected models through, such as voting, weighted averaging, and similar approaches to improve the final decision-making accuracy and stability. Dedicated methods allocate different sub-tasks (e.g., different missing modality cases) to specialized individual models, focusing on specific sub-tasks or sub-datasets. In the discrete scheduler method, users can use natural language instructions to enable the LLMs to autonomously select the appropriate model based on types of modalities and downstream tasks.
(a) Model-Based Ensemble Method
(b) Modality-Based Ensemble Method
Figure 15. Two general architectures for ensemble methods. (a) The model-based ensemble methods are equipped with $n$ three-modal DNN models and obtain the final output based on the prediction results of all $n$ models. (b) In modality-based ensemble methods, a uni-modal DNN is equipped for each modality, and finally the results of these uni-modal DNNs are sent to ensemble methods.
2.4.1. Ensemble Method
@@
Ensemble learning methods allow flexibility in supporting different numbers of expert models to combine their predictions, as shown in Figure 15. Early work (Wagner et al., 2011) in sentiment analysis employed ensemble learning to handle missing modalities, averaging predictions from modality-specific models. With deep learning advancements, the Ensemble-based Missing Modality Reconstruction network (Zeng et al., 2022) was introduced, leveraging weighted judgments from multiple full-modality models when generated missing modality features exhibit semantic inconsistencies. This type of model-based ensemble method, depicted in Figure 15(a), integrates various full-modality models to aid decision-making. Another type is modality-based ensemble methods, as shown in Figure 15(b), where each modality is paired with a uni-modal model, and only available modalities contribute to decision-making. In multimodal medical image diagnosis, early studies found uniformly weighted methods performed better than weighted averaging and voting approaches (Yuan et al., 2012). In multimodal object detection, Chen et al. (Chen et al., 2022) proposed a probabilistic ensemble method that flexibly handles missing modalities via probabilistic marginalization, proving highly efficient. Li et al. (Li et al., 2023a) recently proposed Uni-Modal Ensemble with Missing Modality Adaptation, training models per modality and performing post-fusion training. Other approaches (Miech et al., 2018) calculate feature-based weights to fuse modalities, where weights reflect feature importance for final predictions. A hybrid method, Dynamic Multimodal Fusion (DynMM) (Xue and Marculescu, 2023), uses gating functions to select uni-/multi-modal models dynamically. Lastly, Proximity-based Modality Ensemble (PME) (Cha et al., 2024) addresses negative fusion by employing a cross-attention mechanism with attention bias, adaptively reducing noise in multi-modal decoding.
@@
Figure 16. General idea of Dedicated Methods. If modality 2 in the sample is missing. Dedicated Methods select the sample-suited model from DNN models library for training and testing. In above figure, the three modalities can form a total of seven models with different modality combinations. For easy understanding, we use colored circles in purple rectangles to represent the modalities that the model can handle. In order to adapt to the missing modality 2 (light green circles), Dedicated Methods usually select the DNN model (purple rectangle) that can handle modality 1 (light orange circles) and modality 3 (light blue circles) for training and testing.
2.4.2. Dedicated Method
@@
Dedicated methods assign different tasks to specialized models. We show the general idea of the kind of method in Figure 16. KDNet (Hu et al., 2020) was the first dedicated method proposed to handle different combinations of missing modalities. It treats uni-modality specific models as student models, learning multimodal knowledge from the features and logits of a multimodal teacher model. Those trained uni-modals can be used for different missing modality cases. Following KDNet, ACNet (Wang et al., 2021) also utilizes this distillation method but introduces adversarial co-training, further improving the performance. Lee et al. (Lee et al., 2023b) proposed missing-aware prompts to address missing modality problems based on the prompt learning by using input- and attention- level prompts for each kind of missing modality case, only 1% of the model total parameters needed fine-tuning for downstream tasks. Similarly, some researchers (Reza et al., 2023) equip different adapter layers for each missing modality case to handle missing modality problems.
@@
2.4.3. Discrete Scheduler Method
In discrete scheduler methods (Figure 17), LLMs act as task controllers, determining the execution order of different discrete sub-tasks broken down from the major task/instruction. While the LLM does not directly process multimodal data, it interprets language instructions and orchestrates task execution across uni- and multi-modal modules. This structured yet flexible approach is particularly effective for outputs requiring sequential tasks, enabling the system to handle any number of modalities and naturally addressing missing modality problems. For example, Visual ChatGPT (Wu et al., 2023b) integrates multiple foundation models with ChatGPT to enable interaction through text or/and images, allowing users to pose complex visual questions, provide editing instructions, and receive feedback within a multi-step collaboration framework. HuggingGPT (Shen et al., 2024) is an LLM-driven agent that manages and coordinates various AI models from Hugging Face. It leverages LLMs for task planning, model selection, and summarization to address complex multimodal tasks. ViperGPT (Surís et al., 2023) combines visual and language tasks into modular subprograms, generating and executing Python code for complex visual queries without additional training to achieve effective outputs. There are other similar discrete scheduler approaches, such as AudioGPT (Huang et al., 2024), MM-REACT (Yang et al., 2023a), and LLaVA-Plus (Liu et al., 2023a).
Some aforementioned dedicated and ensemble methods can flexibly handle the missing modality problem without additional training, but most of them require more model storage space, which is not feasible for many resource-constrained devices. For instance, DynMM requires storing various uni- and multi-modality models. As the number of modalities increases, the required number of models also rises. Also, modality-based ensemble methods struggle to adequately model the complex inter-modality relationships to make final predictions. In addition, discrete scheduler methods can solve a variety of tasks when there are sufficient types of downstream modules, but they usually require LLMs to respond quickly and understand human instructions accurately in real world scenarios.
2.5. Discussion
Table 1. Comparison of thirteen kinds of deep multimodal learning with missing modality methods from four dimensions.
Dimension Method Pros and Cons
Modality Augmentation Modality Composition Pros: Simple and effective for data augmentation. Cons: Unsuitable for pixel-level tasks (Composition); Modality number $\uparrow$, generative model number or training complexity $\uparrow$.
@@ 2.5. Discussion
Feature Space Engineering Regularization-based Pros: Flexible for novel modalities with better generalization. Cons: Hard to balance constraints and handle dataset imbalance.
Correlation Driven
@@
Architecture Engineering Attention-based Pros: Attention methods scale well; Distillation and graph learning are effective. Cons: Most attention methods demand high computational resources; Distillation and graph learning struggle with incomplete datasets and training complexity.
@@
Model Selection Ensemble Pros: Effective for specific tasks. Cons: Grows complex with more modalities; Suffers from modality imbalance; Demands significant storage space.
@@ 2.5.1. Summary
In this section, we categorize existing deep MLMM approaches across four dimensions: modality augmentation, feature space engineering, architecture engineering, and model selection. We further subdivide these into thirteen categories under a fine-grained methodology taxonomy. Table 1 summarizes the overall pros and cons of these methods from four dimensions. Generative and distillation methods are the most common approaches for addressing missing modalities; they are easy to implement and deliver strong performance. With the rise of Transformers, attention-based methods have become more popular due to their larger receptive fields and parallelism. However, indirect-to-task generation methods and most distillation methods (except for some intermediate and self-distillation methods) are currently unable to handle incomplete training datasets (means cannot access missing modality data during training). Below, we summarize the advantages and disadvantages of these thirteen methods from four dimensions.
@@
Modality Augmentation: Modality composition methods operate directly on the original data level to address missing modality problems by combining existing data samples or padding missing data. These methods are straightforward and can serve as conventional data augmentation techniques. However, they are typically unsuitable for pixel-level downstream tasks and heavily rely on the existing modalities. On the other hand, modality generation methods employ generative models to synthesize missing modalities, mitigating the performance degradation caused by the missing data. Despite their advantages, these methods are often limited by the availability of full-modality samples in the dataset and the increased training complexity due to the additional modalities. They also require extra storage space for the generative models.
@@
Feature Space Engineering: Regularization- and correlation-based representation constraints allow novel modalities to be introduced by simply adding corresponding branches. However, as the number of constraints increases, balancing them becomes challenging. Moreover, it is difficult to capture unique relationships across all modalities, especially when the dataset is highly imbalanced due to missing modalities. Correlation-based methods, in particular, often fail to handle missing modality samples in the test environment, as they are typically designed for pre-training with incomplete datasets. Similar to modality composition methods, representation composition methods attempt to recovery missing modality representations by combining representations of available modalities. However, since representations typically carry more generalized information than original data, these methods tend to yield better results. Representation generation methods further improve this process by generating representations based on relationships between modalities, allowing the reconstruction of missing modality representations. Many studies have confirmed the effectiveness of this approach in handling missing modalities. However, if the dataset inherently contains missing modalities, indirect-to-task representation generation methods that aim to recovery missing modality representations through the reconstruction of original data are not feasible. Furthermore, when the dataset has severely imbalanced modality combinations, generative methods may also fail, becoming overly reliant on existing modalities.
@@
Architecture Engineering: Some methods focus on designing model training or inference architecture to alleviate the performance degradation caused by missing modalities. Attention-based methods are valued by researchers because they can effectively capture the relationships between various modalities, are scalable in terms of dataset size, and are highly parallelizable. Single attention mechanisms used for modality fusion have the advantage of being plug-and-play. However, the drawbacks of transformer-based methods in this category are also relatively obvious: training multimodal transformers from scratch requires excessive training time and many GPU resources. Recently popular PEL methods can effectively mitigate this drawback, but the generalization performance of this method is still not as well as full-parameter training or tuning. Although MLLMs can handle an arbitrary number of modalities with flexibility, they are constrained by training complexity and require substantial computational resources. Distillation-based methods are relatively easy to implement, with student models learning how to reconstruct missing modality representations and inter-modal relationships from teacher models. Since the teacher model typically receives full-modality samples as input, it can simulate any missing modality cases during training, ensuring strong performance. However, most distillation-based methods are limited to complete datasets and are inapplicable to datasets with inherent missing modalities. To our knowledge, only some intermediate distillation and self-distillation methods have attempted to input mismatched samples of the same class into the teacher for distillation purposes. Graph learning-based methods capture intra- and inter-modal relationships more effectively, but as the number of modalities increases, the development complexity and inefficiency of these methods become more pronounced. They are also currently unsuitable for large-scale datasets.
@@
Model Selection: These methods select one or more models for prediction. Ensemble and dedicated methods are affected by imbalances in modality combinations within the dataset. As the number of modalities grows, the complexity and models to be trained also increase, especially for dedicated methods. The discrete scheduler method requires an LLM to coordinate, but if callable modules are insufficient, downstream tasks may not be completed. Additionally, inference speed is limited by the LLM and modules. Model selection methods also require significant model storage space.
2.5.2. Recovery and Non-Recovery Methods
@@
Table 2. Another common taxonomy of recovery and non-recovery methods based on the three stages of early, intermediate and late in the conventional taxonomy of multimodal learning.
@@
Representation Generation (Ma et al., 2021a)
Distillation-based (Li et al., 2022a)
Correlation Driven (Ma et al., 2021b)
@@
Non-Recovery Dedicated (Wang et al., 2021) Discrete Scheduler (Wu et al., 2023b) Representation Composition (Delgado-Escano et al., 2021) Attention-based (Ge et al., 2023) Graph Learning-based (Zhao et al., 2022) Multimodal Large Language Model (Wu et al., 2023a) Regularization-based (Liang et al., 2019) Distillation-based (Shi et al., 2024) Ensemble (Chen et al., 2022) /
Specifically, we have further classified our method categories in Table 1 into another common taxonomy to facilitate subsequent researchers to better distinguish, which is based on the three-stage (early, intermediate, and late) conventional taxonomy. In Table 2, we provide this taxonomy of our method categories dealing with missing modalities across these stages, based on whether they recover the missing modalities or not. We also attach a reference work for different methods. In our analysis of 315 papers, we found that 75.5% of the works focus on recovering missing modality information for inference, while only 24.6% explore inference without modality recovery. Among the recovery methods, early-stage and mid-stage modality recovery account for 20.3% and 45.8%, respectively, with late-stage and multi-stage recovery methods each accounting for 4.7%. For non-recovery methods, early, mid, and late-stage modality fusion represent 4.2%, 14.1%, and 6.3%, respectively. No non-recovery methods that combine multiple stages were found. We can observe that methods on recovering missing modality features in the middle stage account for the largest proportion. We think this is because recovering features, as opposed to the original data, can avoid much noise and bias while providing more modality-specific information. Additionally, compared to later-stage features, mid-stage features tend to be more enriched.
2.5.3. Limitation
Due to the variety of settings in most missing-modality works, it is difficult to compare the performance of different methods using specific dataset benchmarks. We leave this issue for future work.

### v3

@@ 4.1.4. Multimodal Large Language Models (LLM)
Most MLLMs employ transformer-based modality encoders, such as CLIP, ImageBind (Girdhar et al., 2023), and LanguageBind (Zhu et al., 2023), which encode multimodal inputs into a unified representation space. Examples include BLIP-2 (Li et al., 2023b), which bridges the modality gap using a lightweight Querying Transformer, and LLaVA (Liu et al., 2024a), which enhances visual-language understanding with GPT-4-generated instruction data (Bubeck et al., 2023). These models are optimized for tasks like Visual Question Answering, Dialogue, and Captioning. Recent advancements extend output generation to multiple modalities, such as images. AnyGPT (Zhan et al., 2024) and NExT-GPT (Wu et al., 2023a) unify modalities like text, speech, and images using discrete representations and multimodal projection adaptors, enabling seamless multimodal interaction. CoDi (Tang et al., 2024) introduces Composable Multimodal Conditioning, allowing arbitrary modality generation through weighted feature summation. There are also some other methods that discard the modality encoders and use linear transformation directly, such as Fuyu (Bavishi et al., 2023) and OtterHD (Li et al., 2023c). Although MLLMs can flexibly handle any number of modalities, they have many disadvantages, such as the inconsistent multi-modal positional encoding, training difficulty, and high GPU resource requirements. Additionally, no specific MLLM benchmarks on missing modality problems have been proposed.
4.2. Model Combinations
@@
Model combinations aim to use selected models for downstream tasks. These methods can be categorized into ensemble, dedicated training, and discrete scheduler methods. Ensemble methods combine predictions from multiple selected models through different types of aggregation methods, such as voting, weighted averaging, and similar approaches to improve the accuracy and stability. Dedicated training methods allocate different sub-tasks (e.g., different missing modality cases) to specialized individual models, focusing on specific sub-tasks or sub-datasets. In the discrete scheduler methods, users can use natural language instructions to enable the LLMs to autonomously select the appropriate model based on types of modalities and downstream tasks. We provide more details about these methods below.
(a) Multimodal Model Ensemble Methods
(b) Unimodal Model Ensemble Methods
Figure 15. Two general architectures for ensemble methods. (a) The multimodal model ensemble methods contain $n$ three-modal DNNs and produce final predictions based on the outputs of all $n$ DNNs. (b) Each modality is processed by a unimodal DNN in unimodal model ensemble methods, where final predictions are produced by aggregating the outputs of all accessible unimodal DNNs.
4.2.1. Ensemble Methods
@@
As detailed in the next paragraphs, ensemble learning methods allow flexibility in supporting different numbers of expert models to combine their predictions, as shown in Figure 15. Multimodal Model Ensemble Methods: Early work (Wagner et al., 2011) in sentiment analysis employed ensemble learning to handle missing modalities, averaging predictions from uni-modal models. With deep learning advancements, the Ensemble-based Missing Modality Reconstruction network (Zeng et al., 2022) was introduced, leveraging weighted judgments from multiple full-modality models when generated missing modality features exhibit semantic inconsistencies. This type of multimodal model ensemble method, depicted in Figure 15(a), integrates various full-modality models to aid decision-making.
Unimodal Model Ensemble Methods: The general architecture of this method is shown in Figure 15(b), where each modality is processed by a uni-modal model, and only the available modalities contribute to decision-making. In multimodal medical image diagnosis, early studies found uniformly weighted methods performed better than weighted averaging and voting approaches (Yuan et al., 2012). In multimodal object detection, Chen et al. (Chen et al., 2022) proposed a probabilistic ensemble method. This method does not require training and can flexibly handle missing modalities through probabilistic marginalization, demonstrating high efficiency in experiments. Li et al. (Li et al., 2023a) recently proposed Uni-Modal Ensemble with Missing Modality Adaptation, training models per modality and performing late-fusion training. Other approaches (Miech et al., 2018) calculate feature-based weights to fuse modalities, where weights reflect feature importance for final predictions.
Hybird Methods: Dynamic Multimodal Fusion (DynMM) (Xue and Marculescu, 2023) uses gating mechanisms to select uni/multi-modal models dynamically. Recently, Cha et al. (Cha et al., 2024) proposed Proximity-based Modality Ensemble (PME) to use a cross-attention mechanism with an attention bias to integrate box predictions from different modalities. PME can adaptively combine box features from uni-/multi-model models and reduce noise in multi-modal decoding.
@@
Figure 16. General idea of dedicated training methods. Assuming that modality 2 is missing, dedicated methods select the sample-suited model from DNN models library for training and testing. In the figure, the three modalities can form a total of seven models with different modality combinations. For easy understanding, we use colored circles in purple rectangles to represent the modalities that the model can handle. In order to adapt to the missing modality 2 (light green circles), dedicated methods usually select the DNN model (purple rectangle) that can handle modality 1 (light orange circles) and modality 3 (light blue circles) for training and testing.
4.2.2. Dedicated Training Methods
@@
Dedicated training methods assign different tasks to specialized models. We show the general idea of those methods in Figure 16. KDNet (Hu et al., 2020) was the first dedicated method proposed to handle different combinations of missing modalities. It treats uni-modality specific models as student models, learning multimodal knowledge from the features and logits of a multimodal teacher model. Those trained uni-modals can be used for different missing modality cases. Following KDNet, ACNet (Wang et al., 2021) also utilizes this distillation method but introduces adversarial co-training, further improving the performance. Lee et al. (Lee et al., 2023b) proposed missing-modality-aware prompts to address missing modality problems based on the prompt learning by using input- and attention- level prompts for each kind of missing modality case. This method only needed 1% of the model parameters to be fine-tuned for downstream tasks. Similarly, some researchers (Reza et al., 2023) introduce different adapter layers for each missing modality case.
@@
4.2.3. Discrete Scheduler Methods
In discrete scheduler methods (Figure 17), LLMs act as controllers, determining the execution order of different discrete steps broken down from the major task/instruction. While the LLM does not directly process multimodal data, it interprets language instructions and orchestrates task execution across uni- and multi-modal modules. This structured yet flexible approach is particularly effective for outputs requiring sequential tasks, enabling the system to handle any number of modalities and naturally addressing missing modality problems. For example, Visual ChatGPT (Wu et al., 2023b) integrates multiple foundation models with ChatGPT to enable interaction through text or images, allowing users to pose complex visual questions, provide editing instructions, and receive feedback within a multi-step collaboration framework. HuggingGPT (Shen et al., 2024) is an LLM-driven agent that manages and coordinates various AI models from Hugging Face. It leverages LLMs for task planning, model selection, and summarization to address complex multimodal tasks. ViperGPT (Surís et al., 2023) combines visual and language tasks into modular subprograms, generating and executing Python code for complex visual queries without additional training to achieve effective outputs. There are other similar discrete scheduler approaches, such as MM-REACT (Yang et al., 2023a) and LLaVA-Plus (Liu et al., 2023a).
Some aforementioned dedicated and ensemble methods can flexibly handle the missing modality problem without additional training, but most of them require more model storage space, which is not feasible for many resource-constrained devices. For instance, DynMM requires storing various uni- and multi-modality models. As the number of modalities increases, the required number of models also rises. Also, unimodal model ensemble methods struggle to adequately model the complex inter-modality relationships to make final predictions. Although the some dedicated fine-tuning methods do not require too much time and consumption of resources such as GPU, it is still difficult to compare with full parameter training. In addition, discrete scheduler methods can solve a variety of tasks when there are sufficient types of downstream modules, but they usually require LLMs to respond quickly and understand human instructions accurately in real world scenarios.
5. Methodology Discussion
Table 1. Comparison of deep multimodal learning with missing modality methods from four types and two aspects.
Aspect Type Method Pros and Cons
Data Processing Modality Imputation Modality Composition Pros: Simple and effective for data augmentation. Cons: Unsuitable for pixel-level tasks (Composition); Modality number $\uparrow$, generative model number or training complexity $\uparrow$.
@@ 5. Methodology Discussion
Representation Focused Coordinated-Representation Pros: Flexible for novel modalities with better generalization. Cons: Hard to balance constraints and handle dataset imbalance.
@@
Strategy Design Architecture Focused Attention-based Pros: Attention methods scale well; Distillation and graph learning are effective. Cons: Most attention methods demand high computational resources; Distillation and graph learning struggle with incomplete datasets and training complexity.
@@
Model Combinations Ensemble Pros: Effective for specific tasks. Cons: Grows complex with more modalities; Suffers from modality imbalance; Demands significant storage space.
@@
In the above Section 3 and 4, we divide the existing MLMM methods into four types from the aspects of data processing and strategy design: modality imputation, representation-focused, architecture-focused, and model combinations. We also further subdivide these four types into twelve categories, exploring a fine-grained methodology taxonomy proposed by us. Table 1 summarizes the overall pros and cons of these methods. Generative and distillation methods are the most common approaches; they are easy to implement and deliver strong performance. With the rise of Transformers, Transformers methods have become more popular due to their larger receptive fields and parallelism. However, indirect-to-task generation methods and most distillation methods (except for some intermediate (Zhang et al., 2023c) and self-distillation methods (Wang et al., 2023a; Shi et al., 2024)) are currently unable to handle incomplete training datasets (means cannot access missing modality data during training). Below we provide a concise analysis of these twelve methods from two aspects and four types.
5.1. Data Processing Aspect
@@
(1) Modality Imputation: Modality composition methods operate directly on the input modality data level to address missing modality problems by combining existing data samples or filling missing data. However, they are typically not good at pixel-level downstream tasks and heavily rely on available modalities. On the other hand, modality generation methods employ generative models to synthesize missing modalities, mitigating the performance degradation caused by missing data. But these methods are often limited by the availability and number of full-modality samples and the increased training complexity due to the additional modalities. They also require extra storage for generative models.
@@
(2) Representation Focused: Coordinated-representation methods allow novel modalities to be introduced by simply adding corresponding branches. However, as the number of constraints and modalities increases, balancing them becomes challenging. Correlation-driven methods often fail to handle missing modality samples in the test environment, as they are typically designed for pre-training with incomplete datasets. Similar to modality composition methods, representation composition methods attempt to recover missing modality representations by combining available-modality representations. Since representations typically carry more generalized information than modality data, these methods tend to yield better results. Representation generation methods further improve this process by generating representations based on relationships between modalities, allowing reconstruction of missing modality representations. Many studies have confirmed the effectiveness of this approach in handling missing modalities. However, if the dataset contains missing modality samples, indirect-to-task representation generation methods that aim to recover missing modality representations through reconstruction of modality data are not feasible. When the dataset has severely imbalanced modality combinations, generative methods may also fail, becoming overly reliant on existing modalities.
5.2. Strategy Design Aspect
@@
(1) Architecture Focused: Some methods focus on designing model training or inference architecture to alleviate the performance degradation caused by missing modalities. Attention-based methods are valued by researchers because they can effectively capture the relationships between various modalities, are scalable in terms of dataset size, and are highly parallelizable. Single attention mechanisms used for modality fusion have the advantage of being plug-and-play. The drawbacks of transformer-based methods in this category are also relatively obvious: training multimodal transformers from scratch requires excessive training time and many GPU resources. Recently popular PEL methods can effectively mitigate this drawback, but the generalization performance of this method is still not as well as full-parameter training or tuning. Although MLLMs can handle an arbitrary number of modalities with flexibility, they are constrained by training complexity and require substantial computational resources.
Distillation-based methods are relatively easy to implement, with student models learning how to reconstruct missing modality representations and inter-modal relationships from teacher models. Since the teacher model typically receives full-modality samples as input, it can simulate any missing modality cases during training, ensuring strong performance. However, most distillation-based methods are limited to the requirement of complete datasets and are inapplicable to datasets with inherent missing modalities. To our knowledge, only one intermediate distillation method (Zhang et al., 2023c) has attempted to input mismatched samples of the same class into the teacher and some self-distillation methods (Wang et al., 2023a; Shi et al., 2024) use representations of available modality branches for distillation purposes.
Graph learning-based methods capture intra- and inter-modal relationships more effectively, but as the number of modalities increases, the development complexity and inefficiency of these methods become more pronounced. They are also currently unsuitable for large-scale datasets.
@@
(2) Model Combinations: These methods select models for prediction. Ensemble and dedicated training methods can be affected by imbalances in modality combinations within the dataset. As the number of modalities grows, the complexity and models to be trained also increase, especially for dedicated methods. Discrete scheduler methods requires an LLM to coordinate, but if callable modules are insufficient, downstream tasks may not be completed. Additionally, inference speed is limited by LLMs and modules. Model combinations methods also require significant model storage.
@@
Table 2. Another common taxonomy of recovery and non-recovery methods in multimodal learning is based on three stages: early, intermediate, and late stage, as seen in the conventional taxonomy. In this context, recovery and non-recovery methods refer to how missing modalities are handled at each stage.
@@
Representation Generation (Ma et al., 2021b)
Distillation-based (Li et al., 2022b)
Coordinated-Representation (Ma et al., 2021c)
@@
Non-Recovery Dedicated (Wang et al., 2021) Discrete Scheduler (Wu et al., 2023b) Representation Composition (Delgado-Escano et al., 2021) Attention-based (Ge et al., 2023) Graph Learning-based (Zhao et al., 2022) Multimodal Large Language Model (Wu et al., 2023a) Coordinated-Representation (Liang et al., 2019) Distillation-based (Shi et al., 2024) Ensemble (Chen et al., 2022) Not Applicable
5.3. Recovery and Non-Recovery Methods & Some Statistics
We have further classified MLMMs using the taxonomy in Table 1 to facilitate researchers to better distinguish MLMMs. This classification is based on the three-stage (early, intermediate, and late) classic multimodal learning taxonomy. Table 2 leverages this classic taxonomy of MLMMs and specifies whether they recover missing modalities or not. We also list reference papers for different methods. In our analysis of 315 papers, we found that 75.5% of the works focus on recovering missing modality information, while only 24.5% explore inference without modality recovery. Among the recovery methods, early-stage and intermediate-stage modality recovery methods account for 20.3% and 45.8%, respectively, with late-stage and multi-stage recovery methods accounting for 4.7%, each. For non-recovery methods, early, intermediate, and late-stage modality fusion methods represent 4.2%, 14.1%, and 6.3%, respectively. We were not able to find any non-recovery methods that combine multiple stages. We can observe that methods on recovering missing modality features in the intermediate stage account for the largest proportion. We think this is because recovering features, as opposed to the raw data, can avoid much noise and bias while providing more modality-specific/shared information. Additionally, compared to later-stage features, intermediate-stage features tend to be more enriched.
Limitations: Due to the variety of settings in most missing-modality works, it is difficult to compare performances of different methods across datasets. This represents a research gap that we encourage the community to address.